OpenStudy (anonymous):

Keesha has 10 more quarters than dimes which total 11.25. How many coins does she have in quarters and dimes

OpenStudy (anonymous):

let q = number of quarters let d = number of times q = 10 + d and 0.25q + 0.10d = 11.25

OpenStudy (shinalcantara):

let 'x' be the number of dimes 'y' be the number of quarters 0.10x + 0.25y = 11.25 ----equation 1 y = 10 + x ----equation 2 substitute equation 2 to equation 1 0.10x + 0.25(10+x) = 11.25 0.10x + 2.5 + 0.25x = 11.25 find x substitute the value of x to equation 2